Solve:

  • dy/dx - 2y = 5y2
1
Expert's answer
2021-11-26T12:21:53-0500

"\frac{dy}{dx}-2y=5y^2."

"\frac{dy}{5y^2+2y}=dx."

"\int{\frac{dy}{5y^2+2y}}=\int dx."

"\int(\frac{1}{2y}-\frac{5}{2(5y+2)})dy=\int dx."

"\frac{1}{2}\ln{\frac{y}{5y+2}}=x+C."

"\frac{y}{5y+2}=Ce^{2x}."

"y=\frac{2e^{2x}}{5(C-e^{2x})}=\frac{2}{5(Ce^{-2x}-1)}."